1. Let X be random variable with values 4,5,6,7,8 and 9




a. Find the mean and the variance of the sample without replacement of n=2.




b. Find a the sampling distribution of the mean and construct it's histogram




2. An electrical firm manufactures light bulbs that have mean life span of 750 hours with a standard deviation of 30 hours. Assume that the mean life span of manufactured light bulbs is almost normal distributed. If a sample of 25 light bulbs are random selected, find the:




a. Probability that their mean life span is less than 740 hours.




b. Probability that their mean life span is greater than 775 hours .




Draw the normal distribution curve for each item.

1
Expert's answer
2022-02-22T07:33:32-0500

1. We have population values "4,5,6,7,8,9," population size "N=6"  and sample size "n=2."

Thus, the number of possible samples which can be drawn without replacements is "\dbinom{6}{2}=15."

a.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Sample\ values & Sample\ mean\ (\bar{x}) \\ \hline
 4,5 & 4.5 \\
 \hdashline
 4,6 & 5 \\
 \hdashline
 4,7 & 5.5 \\
 \hdashline
 4,8 & 6 \\
 \hdashline
 4,9 & 6.5 \\
 \hdashline
 5,6 & 5.5 \\
 \hdashline
 5,7 & 6 \\
 \hdashline
 5,8 & 6.5 \\
 \hdashline
 5,9 & 7 \\
 \hdashline
 6,7 & 6.5 \\
 \hdashline
 6,8 & 7 \\
 \hdashline
 6,9 & 7.5 \\
 \hdashline
 7,8 & 7.5 \\
 \hdashline
 7,9 & 8 \\
 \hdashline
 8,9 & 8.5 \\
 \hdashline
\end{array}""\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & \bar{x} & f & f(\bar{x}) & \bar{x}f(\bar{x})& \bar{x}^2 f(\bar{x})\\ \hline
 & 4.5 & 1 & 1/15 & 9/30 & 81/60 \\
 \hdashline
 & 5 & 1 & 1/15 & 10/30 & 100/60 \\
 \hdashline
 & 5.5 & 2 & 2/15 & 22/30 & 242/60 \\
 \hdashline
 & 6 & 2 & 2/15 & 24/30 & 288/60 \\
 \hdashline
 & 6.5 & 3 & 3/15 & 39/30 & 507/60 \\
 \hdashline
 & 7 & 2 & 2/15 & 28/30 & 392/60 \\
 \hdashline
 & 7.5 & 2 & 2/15 & 30/30 & 450/60 \\
 \hdashline
 & 8 & 1 & 1/15 & 16/30 & 256/60 \\
 \hdashline
 & 8.5 & 1 & 1/15 & 17/30 & 289/60 \\
 \hdashline
 Sum= & & 15 & 1 & 6.5 & 521/12 \\
 \hdashline
\end{array}"


"\mu_{\bar{x}}=E(\bar{X})=6.5"

"Var(\bar{X})=\sigma_{\bar{x}}^2=E(\bar{X}^2)-(E(\bar{X}))^2"

"=521\/12-(13\/2)^2=7\/6"

b.

"\begin{matrix}
x\ \ \ \ 4.5 & 5 & 5.5 & 6 & 6.5 & 7 & 7.5 & 8 & 8.5\\
\\
 p(x)\ \dfrac{1}{15} & \dfrac{1}{15} & \dfrac{2}{15} & \dfrac{2}{15} & \dfrac{1}{5} &\dfrac{2}{15} & \dfrac{2}{15} & \dfrac{1}{15} & \dfrac{1}{15}
\end{matrix}"




2. Let "X=" the mean life span: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=750 \ h, \sigma=30\ h, n=25."

a.


"P(X<740)=P(Z<\dfrac{740-750}{30/\sqrt{25}})"

"=P(Z<-5/3)\approx0.0478"




b.


"P(X>775)=1-P(X\le 775)"

"=1-P(Z\le\dfrac{775-750}{30/\sqrt{25}})"

"=1-P(Z\le-25/6)\approx0.000015"